Find aba^-1 where (I) a= 5,7,9 b=1,2,3 (ii) a= (1,2,5)(3,4) b=(1,4,5)

1
Expert's answer
2022-01-19T15:19:37-0500

Let us find "aba^{-1}" where

(i) "a=(5,7,9),\ b=(1,2,3)"


It follows that "a^{-1}=(5,9,7)." Taking into account that the circles "a" and "b" are independent, and hence they commute, we conclude that

"aba^{-1}=(5,7,9) (1,2,3)(5,9,7)=(1,2,3)(5,7,9)(5,9,7)=(1,2,3)=b."



(ii) "a=(1,2,5)(3,4),\ b =(1,4,5)"


It follows that "a^{-1}=(3,4)(1,5,2)." Therefore,

"aba^{-1}=((1,2,5)(3,4))(1,4,5)((3,4)(1,5,2))=((134)(2,5))((3,4)(1,5,2))\n=(123)."